OpenStudy (anonymous):

Can you show me how you get from ln(9)/3ln(7) to 0.3764?

OpenStudy (anonymous):

Step by step please.

OpenStudy (anonymous):

you have no choice but to use a calculator

OpenStudy (anonymous):

there is no way to do this by hand i recommend wolfram

OpenStudy (anonymous):

you can alter the way you write the ln as ln (9/7³) 3ln7 = ln 7³ ln a/ ln b = ln (a/b)

OpenStudy (anonymous):

its minus

Miracrown (miracrown):